# Bayesian Stability Selection and Inference on Inclusion Probabilities

### v1

@@ 2.3 Selecting non-informative and informative priors
Here, we propose a method for translating background knowledge into the Beta distributions. As mentioned before, when $\alpha_{j},\beta_{j}\geq 1$, an increase in $\alpha_{j}$ reflects a greater number of selections, thereby shifting the probability distribution towards higher values, whereas an increase in $\beta_{j}$ reflects a greater number of non-selections, shifting the distribution towards lower values. The sum $\gamma_{j}=\alpha_{j}+\beta_{j}$ corresponds to the total number of pseudo-observations considered on the basis of prior information. It is essential that this total does not exceed the number of observed stability selection iterations $B$, to ensure that the prior does not disproportionately influence the inference.
@@
In the absence of prior information, we recommend using $\alpha_{j}=\beta_{j}=1$ as a non-informative prior distribution. As discussed in Section 1, Bayesian variable selection methods integrate background knowledge into the prior distribution by translating expert knowledge into statistical parameters. When such background knowledge is available, we propose a two-step approach to convert this knowledge into the Beta distribution. The first step involves assessing the level of subjectivity included in the final results, while the second step evaluates the perceived importance of the variables under consideration. Specifically, we pose two questions to data owners for each variable subject to selection.
@@
$\frac{\gamma_{j}}{\gamma_{j}+B}=\tilde{\zeta}_{j}.$ (6)
Conceptually, $\lfloor\gamma_{j}\rfloor$ rows are added to $M(\lambda)$ based on the provided prior information.
@@
Suppose that you plan to conduct 100 experiments to evaluate the relevance of the $j$th variable. Based on your expertise, what percentage of these experiments do you expect to indicate that the $j$th variable is significant?
@@
We denote the response to Question 2 by $\tilde{\xi}_{j}$, where the value $\tilde{\xi}_{j}$ represents the perceived significance of the $j$th variable. Specifically, $\tilde{\xi}_{j}$ corresponds to the proportion of 1’s in the binary pseudo-observations of the $j$th variable. Based on this, we can calculate the values of $\alpha_{j}$ and $\beta_{j}$ as follows
@@
$\alpha_{j}=\lfloor\tilde{\xi}_{j}\gamma_{j}\rfloor\quad\text{and}\quad\beta_{j}=\lfloor(100\%-\tilde{\xi}_{j})\gamma_{j}\rfloor.$ (7)
Thus, $\tilde{\xi}_{j}$ elucidates the values of the prior parameters, providing insight into the relevance of the $j$th variable based on prior knowledge provided.
@@
Therefore, Equations (6) and (7) offer an interpretable procedure to incorporate background knowledge into statistical distributions. Equation (6) sets the total number of pseudo-observations $\gamma_{j}$, while Equation (7) determines the value of the shape parameters $\alpha_{j}$ and $\beta_{j}$ according to the perceived importance of variables. This approach effectively integrates expert opinions into the Beta distribution treating $B$ as a shared parameter between the two frequentist and Bayesian paradigms to control the degree to which prior knowledge impacts the final results.
@@
To the best of our knowledge, existing Bayesian variable selection methods typically engage in inquiries regarding variable importance, similar to the second question we proposed. This approach seamlessly integrates expert background knowledge into the selection process, thus enhancing the management of the subjectivity-objectivity trade-off.
@@ 3.1 Synthetic Data
We generate data with a sample size of $n=50$ and the number of predictor variables $p=500$, where the predictor variables $\boldsymbol{x}^{\top}_{i}$ are independently drawn from $\mathcal{N}(0,\Sigma)$ where $\Sigma$ is the identity matrix except for the elements $\Sigma_{12}=\Sigma_{34}=\Sigma_{35}=\Sigma_{45}=0.8$ and their symmetric counterparts; therefore, two sets of covariates are correlated: $C_{1}=\{x_{1},x_{2}\}$ and $C_{2}=\{x_{3},x_{4},x_{5}\}$. The response variable is linearly dependent only on the first six predictor variables with coefficient vector
$\boldsymbol{\beta}=(0.9,0.9,0.7,0.7,0.7,1.5,0,\ldots,0)^{\top}$
and the error term $\boldsymbol{\varepsilon}$ is an i.i.d. sample of the zero-centred Normal noise with standard deviation $\sigma=2$. A roughly similar scenario, in which the predictor variables exhibit correlation, is considered in the discussion section of Meinshausen and Bühlmann (2010), by Kirk et al. They showed that in such cases, LASSO tends to select one of the variables from the correlated group at random. As a result, the selection frequencies either favour one variable with a high selection frequency or result in low selection frequencies for all variables of the group. To address this issue, they suggested the use of elastic net (Zou and Hastie, 2005), which combines the LASSO $l_{1}$-penalty with the ridge regression $l_{2}$-penalty (Tikhonov, 1963; Hoerl and Kennard, 1970). We employ the elastic net in this context and demonstrate that the problem in such a case persists. In fact, this is one of the situations in which incorporating background knowledge can assist the algorithm in overcoming data inherent ambiguity.
@@ 3.2 Riboflavin Data
The first real data set we use is the well-established ‘riboflavin’ data focused on the production of riboflavin (vitamin B2) from different Bacillus subtilis, provided by Dutch State Mines Nutritional Products, which is available through the hdi R package (Dezeure et al., 2015). The data set comprises a single continuous response variable, representing the logarithm of the riboflavin production rate. In addition, it includes $p=4,088$ covariates, which correspond to the logarithm of the expression levels of $4,088$ bacteria genes. The primary objective of analysing these data is to identify genes that influence riboflavin production, with the ultimate aim of genetically engineering the bacteria to enhance the yield of riboflavin. Data were collected from $n=71$ relatively homogeneous samples, which were repeatedly hybridised during a feed batch fermentation process involving different engineered strains and varying fermentation conditions.
@@
Huang et al. (2008) applied adaptive LASSO (Zou, 2006) to analyse this data and identified $19$ gene probes with non-zero effects on the TRIM32 probe. Subsequently, Li et al. (2017) demonstrated that four probes—1389584_at, 1383996_at, 1382452_at, and 1374106_at—are sufficient to model the response variable within a linear framework. With the exception of the third probe, the other probes were also identified by Huang et al. (2008).
@@ 4.1 Synthetic Data
We present the results of applying Bayesian stability selection to synthetic data. We use the elastic net with the mixing parameter $\texttt{alpha}=0.2$. To determine the appropriate $\lambda$ value, we employ a $10$-fold cross-validation using the cv.glmnet function from the glmnet package in R (Friedman et al., 2010) on the entire data set $\mathcal{D}$ and determine the penalty value using the 1-se rule. We generate $100$ different data sets based on the setting described in Section 3 using different random seeds. We set the number of sub-samples $B=100$ and, at the moment, perform the stability selection without using the Bayesian framework.
The average of the selection frequencies of the first six variables that are the relevant variables over the $100$ generated data sets
$\overline{\hat{\Pi}}_{j}=(0.529,0.546,0.604,0.609,0.622,0.540)^{\top}.$
@@
Although the regression coefficient of the variables in $C_{1}$ is higher than those in $C_{2}$, the average selection frequency of the variables in $C_{1}$ is lower than that of $C_{2}$. In addition, although the regression coefficient of $x_{6}$ is double that of the variables in $C_{2}$, its average selection frequency is lower than the variables in $C_{2}$. This suggests that correlated variables, as a group, attract more attention during the selection process compared to uncorrelated variables. In addition, the size of the correlated group directly influences the selection frequencies, with larger groups having higher selection frequencies. Therefore, this simulation reveals that when the relevant variables are highly correlated and the signal-to-noise ratio is small, even the elastic net may fail to address the random selection issue.
@@
Meinshausen and Bühlmann (2010) states that the selection threshold $\pi_{\text{thr}}$ should be selected from the interval $[0.6,0.9]$ therefore, considering $\pi_{\text{thr}}=0.6$, only variables from $C_{2}$ can be considered stable variables. This underscores the importance of incorporating domain knowledge, which can enhance sampling probabilities and help mitigate the ambiguity of the data through prior knowledge and expertise.
@@
Assume that we have selection frequencies, similar to the average selection frequencies mentioned above, that is, $\hat{\Pi}_{j}^{\lambda}=\overline{\hat{\Pi}}_{j}$. Thus, we can say that $n_{j}$ equals 53, 55, 60, 61, 62, and 54 for the first six variables. As discussed in Sections 1 and 2, Bayesian stability selection provides a method for involving data owners’ knowledge in the analysis in a coherent manner. In this example, if we are informed that for these six variables, the domain expert wants half of the final results to be influenced by background knowledge, that is, $\tilde{\zeta_{j}}=50\%$, and these variables are perceived to have an importance of $70\%$, that is, $\tilde{\xi_{j}}=70\%$, using Equations 6 and 7, we derive $\alpha_{j}=70$ and $\beta_{j}=30$. Consequently, applying Equation 5, the mean of the posterior inclusion probabilities for the first six variables $\mathbb{E}[\pi(\Pi_{j}^{\lambda}\mid M(\lambda))]$ is 0.615, 0.625, 0.650, 0.655, 0.660, and 0.620. Therefore, based on the posterior means and adopting $\pi_{\text{thr}}=0.6$, all six relevant variables can be retrieved.

### v2

@@ 2.3 Assigning non-informative and informative priors
Here, we propose a method for translating background knowledge into the Beta distributions. As mentioned before, when $\alpha_{j},\beta_{j}\geq 1$, an increase in $\alpha_{j}$ reflects a greater number of selections, thereby shifting the probability distribution towards higher values, whereas an increase in $\beta_{j}$ reflects a greater number of non-selections, shifting the distribution towards lower values. The sum $\gamma_{j}:=\alpha_{j}+\beta_{j}$ corresponds to the total number of pseudo-observations considered on the basis of prior knowledge. It is essential that this total does not exceed the number of observed stability selection iterations $B$, to ensure that the prior does not disproportionately influence the results. Although Bayesian statistics does not impose restrictions on the use of priors that may dominate the data, our objective is to maintain a balanced interplay between the objectivity of the data and the subjectivity of the priors, thereby ensuring fair decision-making. As a result, if the stability selection outcomes do not support selection of a variable, the prior will not be too strong to select it and vice versa. In fact, we argue that informed variable selection requires a minimum level of support from both data-driven results and prior knowledge. In the absence of prior information, we recommend using $\alpha_{j}=\beta_{j}=1$ as a non-informative prior distribution.
@@
Bayesian variable selection methods integrate background knowledge into the prior distribution by translating expert knowledge into statistical parameters. When such background knowledge is available, we propose a two-step approach to convert this knowledge into the Beta distribution. The first step involves assessing the level of subjectivity included in the final results, while the second step evaluates the perceived relevance of the variables under consideration. Specifically, we pose two questions to data owners for each variable subject to selection.
@@
$\tilde{\zeta}_{j}=\frac{\gamma_{j}}{\gamma_{j}+B}.$ (5)
Conceptually, $\lfloor\gamma_{j}\rfloor$ rows are added to $M(\lambda)_{j}$ based on the provided prior information.
@@
Based on your knowledge and expertise, what percentage of the data-driven experiments do you expect to indicate that the $j$th variable is relevant?
@@
We denote the response to Question 2 by $\tilde{\xi}_{j}$, where the value $\tilde{\xi}_{j}$ represents the perceived relevance of the $j$th variable. Specifically, $\tilde{\xi}_{j}$ corresponds to the proportion of 1’s in the binary pseudo-observations of the $j$th variable. Based on this, we can calculate the values of $\alpha_{j}$ and $\beta_{j}$ as follows
@@
$\alpha_{j}=\lfloor\tilde{\xi}_{j}\gamma_{j}\rfloor\quad\text{and}\quad\beta_{j}=\gamma_{j}-\alpha_{j}.$ (6)
Thus, $\tilde{\xi}_{j}$ determines the values of the prior parameters, providing insight into the relevance of the $j$th variable based on prior knowledge provided. We expect that $\tilde{\xi}_{j}$ will tend to values closer to either $0\%$ or $100\%$, as it is unreasonable that prior expertise would justify assigning a variable an importance of approximately $50\%$. In the absence of prior information, the use of non-informative priors is recommended.
@@
Therefore, Equations (5) and (6) offer an interpretable procedure to incorporate background knowledge into statistical distributions. Equation (5) sets the total number of pseudo-observations $\gamma_{j}$, while Equation (6) determines the value of the shape parameters $\alpha_{j}$ and $\beta_{j}$ according to the perceived relevance of variables. This approach effectively integrates expert opinions into the Beta distribution treating $B$ as a shared parameter between subjective and objective aspects of the procedure to control the degree to which prior knowledge impacts the final results.
@@
To the best of our knowledge, existing Bayesian variable selection methods typically engage in inquiries regarding variable relevance, similar to the second question we proposed. Our approach seamlessly integrates expert’s background knowledge into the selection process, thus enhancing the management of the subjectivity-objectivity trade-off.
In this paper, previously published results are used to construct informative priors; however, the lack of access to domain experts for our data examples limits our ability to directly demonstrate the effectiveness of the proposed approach. These priors serve as a proxy for expert knowledge, thereby allowing us to simulate the incorporation of expert insights into our methodology.
The variance of the posterior selection probability in the non-informative mode is solely dependent on $B$ and $n_{j}$. However, in the informative mode, it is also influenced by the parameters $\alpha_{j}$ and $\beta_{j}$. Here, for visualization purposes, we assume $\gamma_{j}=B$, which implies that $\tilde{\zeta}=50\%$. Figure 1 illustrates the variance for both non-informative and informative modes, with the total number of sub-samples held constant while $n_{j}$ takes three distinct values. The effect of $\alpha_{j}$ is evaluated across its entire admissible domain. As shown in the figure, for all three values of $n_{j}$, the variance in the informative mode is consistently lower than that in the non-informative mode for all values of $\alpha_{j}$. According to Nogueira et al. (2018), stability is inversely related to variance. Therefore, based on the Figure 1, it can be concluded that in the scenario considered, the informative mode consistently leads to more stable decisions.
(a) $n_{j}=30$
(b) $n_{j}=50$
(c) $n_{j}=70$
Figure 1: Variance of the posterior selection probability as a function of $\alpha_{j}$ when $B=100$ and $n_{j}\in\{30,50,70\}$
The dynamics changes for extreme values of $n_{j}$. As shown in Figure 2, for small or large values of $n_{j}$, the variance in the informative mode is lower than that in the non-informative mode only when prior knowledge aligns with the results derived from the data. In cases where there is a significant discrepancy between expert knowledge and data-driven results, the variance in the informative mode exceeds that in the non-informative mode. In such scenarios, it would be valuable to further investigate the underlying reasons for the strong disagreement between expert opinion and data.
(a) $n_{j}=10$
(b) $n_{j}=90$
Figure 2: Variance of the posterior selection probability as a function of $\alpha_{j}$ when $B=100$ and $n_{j}\in\{10,90\}$
Figure 3 illustrates the variance of the posterior selection probability by allowing both $\alpha_{j}$ and $n_{j}$ to take all admissible values. As shown in Figure 3a, in the non-informative prior case, the variance remains independent of $\alpha_{j}$, while $n_{j}$ exhibits a quadratic effect on the variance. The highest variance occurs when $n_{j}$ is at the midpoint of its range. On the other hand, Figure 3b shows that, for informative priors, maximum variance occurs when $\alpha_{j}=B-n_{j}$, indicating that prior knowledge actively opposes the selection of variables based on data-driven results. In Figure 3b, the minimum variance is observed when both $\alpha_{j}$ and $n_{j}$ are strongly aligned, either in favor or against the selection of the variable.
(a) Non-informative prior
(b) Informative prior
Figure 3: Variance of the posterior selection probability as a function of $\alpha_{j}$ and $n_{j}$ when $B=100$
@@ 3.1 Synthetic Data
In the first scenario, we generate data with a sample size of $n=50$ and the number of predictor variables $p=500$, where the predictor variables $\boldsymbol{x}^{\top}_{i}$ are independently drawn from $\mathcal{N}(0,\Sigma)$ where $\Sigma$ is the identity matrix except for the elements $\Sigma_{12}=\Sigma_{34}=\Sigma_{35}=\Sigma_{45}=0.8$ and their symmetric counterparts; therefore, two sets of covariates are correlated: $C_{1}=\{x_{1},x_{2}\}$ and $C_{2}=\{x_{3},x_{4},x_{5}\}$. The response variable is linearly dependent only on the first six predictor variables with coefficient vector $\boldsymbol{\beta}=(0.9,0.9,0.7,0.7,0.7,1.5,0,\ldots,0)^{\top}$ and the error term $\boldsymbol{\varepsilon}$ is an i.i.d. sample of the zero-centered Gaussian distribution with standard deviation $\sigma=2$. A roughly similar scenario, in which the predictor variables exhibit correlation, is considered in the discussion section of Meinshausen and Bühlmann (2010), by Kirk et al. They showed that in such cases, Lasso tends to select one of the variables from the correlated group at random. As a result, the selection frequencies either favor one variable with a high selection frequency or result in low selection frequencies for all variables of the group. To address this issue, they suggested the use of elastic net (Zou and Hastie, 2005), which combines the Lasso $\ell_{1}$-penalty with the ridge regression $\ell_{2}$-penalty (Tikhonov, 1963; Hoerl and Kennard, 1970). We employ the elastic net in this context and demonstrate that the problem in such a case persists. In fact, this is one of the situations in which incorporating background knowledge can assist the algorithm in overcoming data inherent ambiguity.
In the second scenario, we change the structure of $\Sigma$ and $\boldsymbol{\beta}$. Here, $\Sigma$ is assumed to follow a decaying structure, defined as $\Sigma_{ij}=0.9^{|i-j|}$. The coefficient vector is specified as $\boldsymbol{\beta}=(0.5,0.4,0.3,0.2,0,\ldots,0)^{\top}$. As in the first scenario, we illustrate how Bayesian stability selection can enhance the decision-making process regarding variable relevance when prior information is available.
@@ 3.2 Riboflavin Data
The first real dataset we use is the well-established ‘Riboflavin’ data focused on the production of riboflavin (vitamin B2) from different Bacillus subtilis, provided by Dutch State Mines Nutritional Products, which is available through the hdi R package (Dezeure et al., 2015). The dataset comprises a single continuous response variable, representing the logarithm of the riboflavin production rate. In addition, it includes $p=4,088$ covariates, which correspond to the logarithm of the expression levels of $4,088$ bacteria genes. The primary objective of analyzing these data is to identify genes that influence riboflavin production, with the ultimate aim of genetically engineering the bacteria to enhance the yield of riboflavin. Data were collected from $n=71$ relatively homogeneous samples, which were repeatedly hybridized during a feed batch fermentation process involving different engineered strains and varying fermentation conditions.
@@
Huang et al. (2008) applied adaptive Lasso (Zou, 2006) to analyze this dataset and identified $19$ gene probes with non-zero effects on the TRIM32 probe. Later, Li et al. (2017) demonstrated that four probes—1389584_at, 1383996_at, 1382452_at, and 1374106_at—are sufficient to model the response variable within a linear framework. With the exception of the third probe, the other probes were also identified by Huang et al. (2008).
@@ 4.1 Synthetic Data
For the first scenario, we use the elastic net with the mixing parameter $\texttt{alpha}=0.2$. To determine the appropriate $\lambda$ value, we employ a $10$-fold cross-validation using the cv.glmnet function from the glmnet package in R (Friedman et al., 2010) on the entire dataset $\mathcal{D}$ and determine the regularization value using the 1-se rule. We generate $100$ different datasets based on the setting described in Section 3 using different random seeds. We set the number of sub-samples $B=100$ and, at the moment, perform the stability selection without using the Bayesian methodology.
@@
The average of the selection frequencies of the first six variables that are the relevant variables over the $100$ generated datasets $\overline{f_{j}^{\lambda}}=(0.529,0.546,0.604,0.609,0.622,0.540)^{\top}$. The maximum selection frequency among irrelevant covariates is $0.062$. Although the regression coefficient of the variables in $C_{1}$ is higher than those in $C_{2}$, the average selection frequency of the variables in $C_{1}$ is lower than that of $C_{2}$. In addition, although the regression coefficient of $x_{6}$ is more than double that of the variables in $C_{2}$, its average selection frequency is lower than the variables in $C_{2}$. This suggests that when using the elastic net, correlated variables, as a group, attract more attention during the selection process compared to uncorrelated variables. In addition, the size of the correlated group directly influences the selection frequencies, with larger groups having higher selection frequencies.
@@
Meinshausen and Bühlmann (2010) stated that the selection threshold $\pi_{\text{thr}}$ should be selected from the interval $[0.6,0.9]$. This choice is reasonable, as for values near $0.5$, the selection frequency does not exhibit a stable dominance between the selection and non-selection of the variable. In this paper, we adopt a threshold of $0.6$, which represents the most conservative value. Therefore, considering $\pi_{\text{thr}}=0.6$, only variables from $C_{2}$ can be considered stable variables. This underscores the importance of incorporating domain knowledge, which can potentially enhance selection frequencies and help mitigate the ambiguity of the data through prior knowledge and expertise.
@@
Assume that for the first six variables, we have selection frequencies, similar to the average selection frequencies mentioned above, that is, $f_{j}^{\lambda}=\overline{f_{j}^{\lambda}}$. Thus, we can say that $n_{j}$ equals 53, 55, 60, 61, 62, and 54 for the first six variables. As discussed in Sections 1 and 2, Bayesian stability selection provides a method for involving data owners’ knowledge in the analysis in a coherent manner. In this example, if we are informed that for these six variables, the domain expert wants half of the final results to be influenced by background knowledge, that is, $\tilde{\zeta_{j}}=50\%$, and these variables are perceived to have a relevance of $70\%$, that is, $\tilde{\xi_{j}}=70\%$, using Equations (5) and (6), we derive $\alpha_{j}=70$ and $\beta_{j}=30$ for $j\in\{1,2,\dots 6\}$. Consequently, the mean of the posterior selection probabilities for the first six variables $\mathbb{E}(\Pi_{j}^{\lambda}\mid n_{j},K)$ is 0.615, 0.625, 0.650, 0.655, 0.660, and 0.620. Therefore, based on the posterior means and adopting $\pi_{\text{thr}}=0.6$, all six relevant variables can be retrieved.
(a) Relevant variables
(b) Irrelevant variables
Figure 4: Heatmap of correct and incorrect selections versus $\tilde{\zeta}_{j}$ and $\tilde{\xi}_{j}$ in the first scenario
Figure 4 presents the heat map that depicts correct and incorrect selections in the first scenario. Figure 4a focuses specifically on the first six variables, where it is assumed that the expert’s responses to the two proposed questions are the same for these variables. This figure shows the number of variables correctly selected under various combinations of $\tilde{\zeta}_{j}$ and $\tilde{\xi}_{j}$. As demonstrated, the optimal outcome occurs when strong prior information regarding the selection of relevant variables is available.
Figure 4b examines the irrelevant variables, totaling 494. Again, it is assumed that the expert’s responses to the two questions are the same for all these variables. The figure shows that in all combinations of $\tilde{\zeta}_{j}$ and $\tilde{\xi}_{j}$, incorrect prior information suggesting the selection of irrelevant variables does not affect the final decision of the method. This favorable outcome arises from design of the method, which ensures that prior information and data-driven results do not disproportionately influence the final selection decisions.
Following a similar approach to the first scenario, and assuming $f_{j}^{\lambda}=\overline{f_{j}^{\lambda}}$, that is, considering the average of selection frequencies for the first four variables across $100$ datasets as the obtained selection frequencies, the selection frequencies for the second scenario are computed as $0.544$, $0.563$, $0.527$, and $0.43$. This implies that none of these variables can be selected. The largest selection frequency among the irrelevant variables is $0.301$. However, the expected posterior selection probabilities $\mathbb{E}(\Pi_{j}^{\lambda}\mid n_{j},K)$ for the first four variables, assuming $\tilde{\zeta}=50\%$ and $\tilde{\xi}=70\%$, are $0.62$, $0.63$, $0.615$, and $0.565$. Consequently, the first three variables can be retrieved.
(a) Relevant variables
(b) Irrelevant variables
Figure 5: Heatmap of correct and incorrect selections versus $\tilde{\zeta}_{j}$ and $\tilde{\xi}_{j}$ in the second scenario
Figure 5 presents the heat map that depicts correct and incorrect selections in the second scenario, similar to Figure 4. Figure 5a focuses specifically on the first four variables. As demonstrated, the optimal outcome is achieved when strong prior information about the selection of relevant variables is available.
Figure 5b examines the irrelevant variables, totaling 496. The figure shows that in almost all combinations of $\tilde{\zeta}_{j}$ and $\tilde{\xi}_{j}$, incorrect prior information suggesting the selection of irrelevant variables does not affect the final decision of the method. Specifically, in the most extreme case, with $\tilde{\zeta}_{j}=50\%$ and $\tilde{\xi}_{j}=100\%$, only two irrelevant variables are incorrectly selected.
It should be stated that $\tilde{\zeta}_{j}$ and $\tilde{\xi}_{j}$ must be carefully calibrated by domain experts, as their appropriate values can vary across different variables under selection depending on the state of knowledge.